OpenStudy (anonymous):

solve by using double angle or half angle formula 1-cos4theta / sin4theta

OpenStudy (anonymous):

is that $$1-\frac{\cos4\theta}{\sin4\theta}$$?

OpenStudy (anonymous):

i think it should be\[\frac{1-\cos 4\theta}{\sin 4\theta}\]

OpenStudy (anonymous):

ok lets start with this:\[\cos4\theta=1-2\sin ^{2}2\theta\] \[\rightarrow 1-\cos4\theta=2\sin ^{2}2\theta\] \[\sin4\theta=2\sin2\theta \cos2\theta\] so then the equation will be:\[2\sin ^{2}2\theta \div 2\sin 2\theta \cos2\theta\] which will give: \[\sin2\theta \div \cos2\theta\] which is :\[\tan2\theta\] i think this is enough simplification..

OpenStudy (anonymous):

thanks so much!